Alex dumped out his piggy bank and there were only nickels, dimes, and quarters.
There were three times as many dimes as nickels, and 40 more quarters than nickels.
If the total amount of money was $17.20, how many of each type of coin were there?

Found 2 solutions by Boreal, ewatrrr:
Answer by Boreal(15235)

nickels=N
dimes=3N
quarters=N+40
.05N+.10(3N)+.25(N+40)=17.20
.60N+10.00=17.20
.60N=7.20
N=12 nickels or $0.60
Dimes=36 or $3.60
quarters=N+40 or 52 or $13.00

Answer by ewatrrr(24785)


Hi
There were three times as many dimes as nickels, and 40 more quarters than nickels (x). 
let x  represent the number of Nickels

  10¢(3x) + 25¢(x+ 40) + 5¢(x) = 1720¢   Cents Makes Sense
   60x = 720
     x = 12  the number of Nickels, 36 Dimes  &  52 Quarters

  5(12) + 10(36) + 25(52) = 1720 checks

Wish You the Best in your Studies.